In four years Ben will be three times as old as he was 12 years ago. How old is he now?
Answer by lwsshak3(11628)

In four years Ben will be three times as old as he was 12 years ago. How old is he now?
**
let x=Ben's age now
x+4 =Ben's age in 4 years
x-12=Ben's age 12 years ago
..
x+4=3(x-12)
x+4=3x-36
2x=40
x=20
ans:
Ben's age now=20